A Paddleboat can move at the speed of 8km/hr in still water. The boat is paddled 10km downstream in the same time it takes to go 5km upstream. What is the speed of the river
Answer by mananth(16946)

Boat speed = 8 km/h
current speed = x km/h

speed against current 8 - x
speed with current 8 + x

Distance against current 5 km
Distance with current 10 km

t=d/r time against = time with current
5/(8-x)=10/(8+x)
5(8+x)= 10(8-x)
40+5x=80-10x
10x+5x =80-40
15x=40
/ 15 check
x = 2.67 km/h current speed